OpenStudy (anonymous):

plz help me to solve this que...... 3e^x tany + (1-e^x)sec^2y dy/dx =0

OpenStudy (anonymous):

\[\frac{3e^x}{e^x-1}dx = \frac{\sec^2y}{\tan y}dy\]Now integrate it, you got 4800 AIR. I expect you to do it.

OpenStudy (anonymous):

ya thnxx budy

OpenStudy (anonymous):

I believe it's 3e^x..?

OpenStudy (anonymous):

Probably wrong though

OpenStudy (anonymous):

ya